OpenStudy (anonymous):

32 - 6x + 7x = -15 + 28

OpenStudy (gabebae):

Are you solving for x?

OpenStudy (anonymous):

yes i should explain better

OpenStudy (gabebae):

Combine all the like terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

32 - 13x = -15 + 28

OpenStudy (anonymous):

im guessing ur afk

OpenStudy (anonymous):

HELLLLLP ANYONEEEEE

rebeccaxhawaii (rebeccaxhawaii):

i gotchu what do you not understand

OpenStudy (anonymous):

just the next step

rebeccaxhawaii (rebeccaxhawaii):

okay well

rebeccaxhawaii (rebeccaxhawaii):

ill break it down for you

rebeccaxhawaii (rebeccaxhawaii):

what you need to do is add like terms

OpenStudy (anonymous):

ok

rebeccaxhawaii (rebeccaxhawaii):

and what you need to do is get the x's on one side and regular numbers on the other

rebeccaxhawaii (rebeccaxhawaii):

32 - 6x + 7x = -15 + 28 but what you need to know is that when you move it to the other side you need to switch the sign so if it was a negative it becomes a positive on the other side

rebeccaxhawaii (rebeccaxhawaii):

so -6x+7x=-15+28-32

rebeccaxhawaii (rebeccaxhawaii):

thats your new problem so solve both sides what do you get ?

OpenStudy (anonymous):

1x = 16

OpenStudy (anonymous):

so x=16

rebeccaxhawaii (rebeccaxhawaii):

there you go

rebeccaxhawaii (rebeccaxhawaii):

wait no

rebeccaxhawaii (rebeccaxhawaii):

you missed something

OpenStudy (anonymous):

thx madam

OpenStudy (anonymous):

huh wut did i miss

rebeccaxhawaii (rebeccaxhawaii):

-15+28-32 is not 16 its -19

rebeccaxhawaii (rebeccaxhawaii):

so x = -19

OpenStudy (anonymous):

lol i forgot the sign thx

OpenStudy (anonymous):

now i can get into 9th grade

OpenStudy (anonymous):

thx

rebeccaxhawaii (rebeccaxhawaii):

hah good job